### v1

# Precise Interception Flight Targets by Image-based Visual Servoing of Multicopter

Abstract
Interception of low-altitude intruding targets with low-cost drones equipped strapdown camera presents a competitive option. However, the malicious maneuvers by the non-cooperative target and the coupling of the camera make the task challenging. To solve this problem, an Image-Based Visual Servoing (IBVS) control algorithm based on proportional navigation guidance with field-of-view holding capability is designed. The proposed controller reduces the miss distance while improving the stability of the visual servo system during interception. Software-in-the-loop (SITL) simulation experiments show a 72.8% reduction in the circular error probability (CEP) compared to the most recent study. This improvement enhances interception accuracy from the decimeter to the centimeter level. Real-world experiments further validate the effectiveness of the proposed algorithm.
Index Terms:
precise interception, IBVS, proportional navigation guidance, multicopter control
I Introduction
THE presence of non-cooperative targets at low altitudes poses a significant threat to flight safety and hinders the development of low-altitude economies[1, 2, 3, 4]. Commonly used countermeasures such as radio-frequency (RF) signal jamming[5, 6] and high-energy weapon shoot-downs[7, 8] are effective but have detrimental environmental impacts. The capture methods[9, 10] are environmentally friendly but less effective against moving targets. All these methods exhibit notable deficiencies when dealing with moving targets. The use of sensor-equipped unmanned aerial vehicles (UAVs) to intercept intruding targets has gained significant attention due to their rapid deployment, safety, and cost-effectiveness. Camera-based solutions, in particular, have shown great potential due to their low cost, lightweight, and high versatility[11, 12, 13].
Figure 1: Intercepting flight target with multicopter.
In recent years, much research has focused on camera-equipped multicopters for intercepting non-cooperative targets. These methods are primarily divided into two categories: Location estimation-based and IBVS-basd. Common Location estimation-based methods employ binocular cameras[11, 14], the Kalman filter[13, 15], and geometry-based approaches[16] for target estimation. After that, a trajectory is planned for interception. However, these methods are error-sensitive and limited by sensor perception range and computational performance. Yet the IBVS methods have a simple structure that calculates the control quantity directly from the sensing module. So it offers advantages such as insensitivity to modeling and calibration, fast response, and extended sensing capabilities. These advantages make it a highly competitive solution for interception tasks.
Multicopter IBVS faces significant challenges in precise interception of non-cooperative flight targets. During interception, the observable field of view (FOV) is susceptible to the multicopter’s attitude, making it challenging to maintain the 2D visibility required for IBVS [17]. Interception accuracy is further affected by target maneuvers, image processing delay, multicopter dynamics delay, and guidance strategy. In our previous work[18, 19], four and six degree of freedom (DOF) IBVS controllers were designed for low and high speed target interception, respectively. Delayed Kalman filtering (DKF) was proposed in [19] for mitigating image processing delay. Some other studies focus on target observation innovations. A virtual-plane IBVS scheme [20] decouples multicopter dynamics and uses an improved image error term to reduce the impact of multicopter rotation during 3D target chasing. A pseudo-linear Kalman filter [21] and a 3D helical guidance law enhances observability for intruding multicopter tracking. Classic guidance laws, such as the pursuit guidance (PG) method [22], aim for lower off-target interception but are less efficient for moving targets. These methods can cause multicopter overload saturation or loss of the target due to image processing latency, inappropriate guidance strategies, and target maneuvers. No related work has focused on the bad impact of overload on interception accuracy. An ideal flat trajectory at the end phase of interception reduces direction angle measurement errors from image processing latency and provides a greater dynamics margin to handle target maneuvers.
In this paper, an IBVS controller is designed for precision interception of non-cooperative flight targets, inspired by bird-catching behaviors. Birds rely solely on vision for efficient hunting, similar to IBVS methods for multicopter interception tasks (see Fig. 1). Specifically, the trajectory of a peregrine falcon hunting mallard duck could described by proportional navigation guidance (PNG) [23], inspired our use of PNG for our tasks. PNG, widely used in missile guidance for decades, flattens the end-phase trajectory and reduces the miss distance[24]. Additionally, a FOV holding controller mimics the bird’s neck rotation for prey tracking. So the proposed controller can continuously and stably track the target and precise interception like a peregrine falcon. Overall, the contributions of this paper are summarized as follows:
  1. 1)
    The proposed algorithm achieved a Circular Error Probability (CEP) of 0.089 meters in typical tests, improving interception accuracy from the decimeter level of recent algorithms [19] to the centimeter level. CEP is a measure where there is a 50% probability that the interceptor will hit within a circle centered on the target with the given radius [25].
  2. 2)
    A FOV holding controller that incorporates PNG and multicopter dynamics is designed to enhance robustness in target tracking.
  3. 3)
    Real and simulated experiments, with extensive comparisons across various scenarios, validate the effectiveness and accuracy of the proposed algorithms.
The rest of this paper is organized into five parts. Section II presents the basic modeling and algorithmic framework. Details of the IBVS controller design are provided in Section III. Simulation experiments are discussed in Section IV, while real-world experiments are covered in Section V. Finally, Section VI concludes with a summary of findings and directions for future research.
II MODELS AND FRAMEWORK
Figure 2: Coordinate system frames and description of the interception problem. In the earth coordinate system, vector $\mathbf{n}_{\mathrm{t}}$ (red vector) represents the LOS direction, while vector $\mathbf{n}_{\mathrm{v}}$ (green vector) represents the velocity direction. This information is obtained from sensors carried by the drone.
II-A Coordinate systems and relative motion models
There are four coordinate systems used, as shown in Figure 2: the Earth-Fixed Coordinate System (EFCS): $o_{e}-x_{\mathrm{e}}y_{e}z_{\mathrm{e}}$, the Body Coordinate System (BCS): $o_{b}-x_{b}y_{b}z_{b}$, Camera Coordinate System (CCS): $o_{c}-x_{c}y_{c}z_{c}$ and Image pixel coordinate system (IPCS): $o_{i}-uv$.
In the relative motion model, both the target ($\mathrm{T}$) and the multicopter ($\mathrm{M}$) are assumed to be point masses. In the EFCS, the relative relationship can be expressed as:
${}^{\mathrm{e}}\mathbf{p}_{\mathrm{r}}$ $={{}^{\mathrm{e}}\mathbf{p}_{\mathrm{T}}}-{{}^{\mathrm{e}}\mathbf{p}_{\mathrm{M}}}$ (1)
${}^{\mathrm{e}}\mathbf{v}_{\mathrm{r}}$ $={{}^{\mathrm{e}}\mathbf{v}_{\mathrm{T}}}-{{}^{\mathrm{e}}\mathbf{v}_{\mathrm{M}}}$
${}^{\mathrm{e}}\mathbf{a}_{\mathrm{r}}$ $={{}^{\mathrm{e}}\mathbf{a}_{\mathrm{T}}}-{{}^{\mathrm{e}}\mathbf{a}_{\mathrm{M}}}$
$={{}^{\mathrm{e}}\mathbf{a}_{\mathrm{r}}^{\mathrm{t}}}+{{}^{\mathrm{e}}\mathbf{a}_{\mathrm{r}}^{\mathrm{n}}}$
where ${}^{\mathrm{e}}\mathbf{a}_{\mathrm{r}}^{\mathrm{t}}$ and ${}^{\mathrm{e}}\mathbf{a}_{\mathrm{r}}^{\mathrm{n}}$ represent the relative tangential and normal acceleration in the EFCS, respectively. The relative position ${}^{\mathrm{e}}\mathbf{p}_{\mathrm{r}}$ also denotes the line-of-sight (LOS) between the interceptor and the target.
II-B Multicopter Dynamics Model
The multicopter used for interception is modeled as a rigid body with mass $m$. Its flight control rigid model can be summarized as [26]:
[…]
where ${}^{\mathrm{e}}\mathbf{p}_{\mathrm{M}}$ and ${}^{\mathrm{e}}\mathbf{v}_{\mathrm{M}}$ is the multicopter’s position and velocity in the EFCS, respectively. $\mathbf{g}$ denotes the acceleration due to gravity, typically represented as $\left[\begin{array}[]{ccc}0&0&g\end{array}\right]^{\mathrm{T}}$ with $g\approx$ 9.8 $m/s^{2}$. ${}^{\mathrm{e}}\mathbf{f}\in\mathbf{R}^{3}$ is the controlled lift of the multicopter, directed opposite to the z-axis of the airframe. $\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\in\mathrm{SO}\left(3\right)$ represents the rotation from the BCS to the EFCS; ${}^{\mathrm{b}}\omega$ is the angular velocity in the BCS, and $\left[{}^{\mathrm{b}}\bm{\omega}\right]_{\times}$ is the cross-product matrix associated with ${}^{\mathrm{b}}\bm{\omega}$, where $\left[\cdot\right]_{\times}$ denotes the matrix such that $\left[\mathbf{x}\right]_{\times}\mathbf{y}=\mathbf{x}\times\mathbf{y}$ for any $\mathbf{x},\mathbf{y}\in\mathbf{R}^{3}$. $\mathbf{J}$ is the moment of inertia of the multicopter, $\mathbf{G}_{\mathrm{a}}$ represents the gyroscopic moment, and ${}^{\mathrm{b}}\mathbf{M}_{\mathrm{d}}$ denotes the aerodynamic moment associated with propeller steering.
II-C Camera Image Model
The multicopter is equipped with a strapdown monocular camera modeled using the pinhole model. This model projects target ${}^{\mathrm{c}}\mathbf{p}_{\mathrm{T}}=\left[{{}^{\mathrm{c}}\mathrm{p}_{tx}}\,\,^{\mathrm{c}}\mathrm{p}_{ty}\,\,^{\mathrm{c}}\mathrm{p}_{tz}\right]^{\mathrm{T}}$ from 3D space to the 2D camera plane: $\mathbbm{R}^{3}\rightarrow\mathbbm{R}^{2}$. The center point of the target in the IPCS is denoted as ${}^{\mathrm{i}}\mathbf{p}=\left[u,v\right]^{\mathrm{T}}$.The error in the IPCS, represented as $\mathbf{e}=\left[e_{x},e_{y}\right]^{\mathrm{T}}$, is defined as:
$\mathbf{e}\triangleq{{}^{\mathrm{i}}\mathbf{p}}-{{}^{\mathrm{i}}\mathbf{p}_{\mathrm{c}}}$ (3)
where ${}^{\mathrm{i}}\mathbf{p}_{\mathrm{c}}=\left[u_{0},v_{0}\right]^{\mathrm{T}}$ is the geometric centroid of the image in the IPCS. The relationship between the error and the change in camera motion can be described by the Jacobian matrix of the IBVS[27]:
$\dot{\mathbf{e}}$ $={\mathbf{L}_{s}}{{}^{\mathrm{c}}\tilde{\mathbf{v}}},$ (4)
$\mathbf{L}_{s}=\left[\begin{matrix}-\frac{1}{{}^{\mathrm{c}}\mathrm{p}_{tz}}&0&\frac{\bar{e}_{x}}{{}^{\mathrm{c}}\mathrm{p}_{tz}}&\bar{e}_{x}\bar{e}_{y}&-\left(1+\bar{e}_{x}^{2}\right)&\bar{e}_{y}\\ 0&-\frac{1}{{}^{\mathrm{c}}\mathrm{p}_{tz}}&\frac{\bar{e}_{y}}{{}^{\mathrm{c}}\mathrm{p}_{tz}}&1+\bar{e}_{y}^{2}&-\bar{e}_{x}\bar{e}_{y}&-\bar{e}_{x}\\ \end{matrix}\right]$
where ${}^{\mathrm{c}}\tilde{\mathbf{v}}=\left[{}^{\mathrm{c}}\mathbf{v}^{\mathrm{T}}\,\,{}^{\mathrm{c}}\mathbf{\omega}^{\mathrm{T}}\right]^{\mathrm{T}}$ reprents the instantaneous linear velocity angular velocity of the camera. ${\bar{\mathbf{e}}}=\left[\bar{e}_{x}\,\,\bar{e}_{y}\right]^{\mathrm{T}}=\left[e_{x}/f_{oc}\,\,e_{y}/f_{oc}\right]^{\mathrm{T}}$ represents the normalized image error of features in IPCS, and $f_{\mathrm{oc}}$ is the focal length of the camera.
The LOS vector direction $\mathbf{n}_{\mathbf{t}}=\left[n_{tx}\,\,n_{ty}\,\,n_{tz}\right]^{\mathrm{T}}$, which connects the target and the multicopter in the EFCS, can then be characterized as:
$\mathbf{n}_{\mathbf{t}}$ $=\frac{{}^{\mathrm{e}}\mathbf{p}_{\mathrm{r}}}{\parallel{{}^{\mathrm{e}}\mathbf{p}_{\mathrm{r}}}\parallel}$ (5)
$\left.=\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\mathbf{R}_{\mathrm{c}}^{\mathrm{b}}\frac{[e_{x}\quad e_{y}\quad f_{\mathrm{oc}}]^{\mathrm{T}}}{\left\|\left[\begin{matrix}e_{x}&e_{y}&f_{\mathrm{oc}}\\ \end{matrix}\right]\right.^{\mathrm{T}}\parallel}\right.$
where $\mathbf{R}_{\mathrm{c}}^{\mathrm{b}}$ is the rotation matrix from the CCS to the BCS. The LOS vector should not exceed the camera’s FOV during interception to fulfill the 2D visibility of the IBVS.
II-D Framework overview
The system framework is shown in Fig. 3. First, the image is processed using a target detection algorithm [28], and corrections are estimated using the DKF [19]. The IBVS controller algorithm then calculates the desired lift $f_{\mathrm{d}}$ and angular velocity $\omega_{\mathrm{d}}$ for the multicopter flight controller. Finally, the controller generates PWM signals for the motors to execute the target interception.
Figure 3: Framework of proposed IBVS controller based on PNG.
III CONTROLLER DESIGN AND ANALYSIS
III-A Guidance Algorithm Design
The PNG calculates the rate of change of velocity proportional to the rate of change of the line of sight (LOS), causing the rate of change of the LOS to become progressively smaller and ultimately achieving interception:
$\frac{\mathrm{d\sigma}}{\mathrm{dt}}=K\frac{\mathrm{dq}}{\mathrm{dt}}$ (6)
where $\sigma$ is the velocity angle, and $q$ is the LOS angle. $K$ is PNG constant. The trajectory before interception becomes progressively flatter and steadily reaches the target when the value exceeds a certain threshold[24]. However, a value too large can make the controller aggressive and sensitive to noise. The value of $K$ is typically chosen between 2 and 6.
The three-dimensional plane is divided into two two-dimensional planes, as shown in Fig. 2. $q_{y}$ and $q_{z}$ are the LOS angles in the vertical and horizontal planes:
$q_{y}=\arctan\left(\frac{n_{tz}}{\sqrt{n_{tx}^{2}+n_{ty}^{2}}}\right);q_{y}\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$ (7)
$q_{z}=\arctan\left(\frac{n_{ty}}{n_{tx}}\right);q_{z}\in(-\pi,\pi).$
And normalized velocity direction vector $\mathbf{n}_{\mathbf{v}}=\left[n_{vx}\,\,n_{vy}\,\,n_{vz}\right]^{\mathrm{T}}$. Thus, $\sigma_{y}$ and $\sigma_{z}$ are the velocity angles in the vertical and horizontal planes which are donated as:
$\sigma_{y}=\arctan\left(\frac{n_{vz}}{\sqrt{n_{vx}^{2}+n_{vy}^{2}}}\right);\sigma_{y}\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$ (8)
$\sigma_{z}=\arctan\left(\frac{n_{vy}}{n_{vx}}\right);\sigma_{z}\in(-\pi,\pi).$
Using the integral Eq. (6), desired velocity angle is calculated from the current moment $k$ and the previous moment $k-1$ as:
$\left\{\begin{array}[]{c}\sigma_{\mathrm{yd}}=K_{y}(q_{y(k)}-q_{y(k-1)})+\sigma_{y(k-1)}\\ \sigma_{\mathrm{zd}}=K_{z}(q_{z(k)}-q_{x(k-1)})+\sigma_{z(k-1)}\\ \end{array}\right.$ (9)
where $K_{y}$ and $K_{z}$ are proportionality coefficient for the vertical and horizontal planes. Timothy J. Layman et al. shows that a value of about 3 is appropriate when multirotor pursuit[29].
Then the desired velocity can be expressed as:
$\mathbf{v}_{\mathrm{d}}=v_{\mathrm{d}}\mathbf{n}_{\mathrm{vd}}$ (10)
where $\mathbf{n}_{\mathrm{vd}}=\left[\sin\sigma_{\mathrm{yd}}\cos\sigma_{\mathrm{zd}}\,\,\sin\sigma_{\mathrm{yd}}\sin\sigma_{\mathrm{zd}}\,\,\cos\sigma_{\mathrm{zd}}\right]^{\mathrm{T}}$. So the desired velocity can be get by designing the desired velocity magnitude.
III-B Field of View Holding Controller
In this subsection, a FOV holding controller is designed to improve the robustness of IBVS during interception. Unlike some control strategies where the z-axis of the CCS is aligned with the target [19] and the target converges to the center of the virtual image plane [20], a new strategy is proposed that aims to:
$\left\{\begin{array}[]{c}e_{x}\rightarrow 0\\ \varDelta e_{y}\leqslant\varepsilon\\ \end{array}\right.$ (11)
Where $\varepsilon$ is a small constant and $\varDelta e_{y}$ is the difference between the largest and smallest $e_{y}$ during the interception. The primary control objective is to make the target converge to the center in the u-axis direction on the image plane. The secondary control objective is to minimize the range of motion in the v-axis direction.
The causes of target movement in the image plane during interception are analyzed. The main factors are the change of ${}^{\mathrm{c}}\mathbf{p}_{\mathrm{r}}$ due to relative motion and the change in attitude $\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}$ of the multicopter. Due to the role of PNG, the LOS change $\varDelta q_{\mathrm{g}}$ is minimal. Therefore, the focus is on improving target tracking and minimizing the adverse effects of attitude changes by designing a suitable FOV holding controller.
Firstly, the relationship between camera motion and $e_{x}$ can represent from Eq. (4) as:
[…]
Due to the decoupling of the yaw angle and dynamics of the multicopter, a PD-based controller is designed for improving target tracking:
${}^{\mathrm{b}}w_{\psi}=k_{\mathrm{p}}e_{x}+k_{\mathrm{d}}\dot{e}_{x}$ (13)
where $k_{\mathrm{p}}$ and $k_{\mathrm{d}}$ are small negative value. This controller improves the settling time and reduces the overshoot. Furthermore, it helps $e_{x}\rightarrow 0$ and is insensitive to pixel noise.
Changes in roll angle at greater distances from the target have less impact on the target in the FOV. Therefore, primary attention is given to changes in the pitch angle. To minimize the adverse effects of attitude changes, an acceleration controller is designed as follows:
${v}_{\mathrm{d}}=v_{\mathrm{now}}+k_{a}$ (14)
where $k_{a}$ is the velocity gain parameter, typically taken as 1-3. Thus, the dynamics-induced LOS error $\varDelta q_{\mathrm{d}}\leqslant\mathrm{arctan}(k_{a}/g)$. The change in the LOS angle of the target in the v direction of the IPCS can be expressed as:
$\varDelta q_{\mathrm{y}}=\varDelta q_{\mathrm{d}}+\varDelta q_{\mathrm{g}}.$ (15)
Based on the geometric relationship, the variation of the target in the image coordinate system can be calculated as:
$\varDelta e_{y}\leqslant\frac{v_{0}\tan\left(\varDelta q_{\mathrm{y}}\right)}{\tan\left(\small{\frac{1}{2}\alpha_{\mathrm{vfov}}}\right)}\leqslant\varepsilon$ (16)
where the parameter $\alpha_{\mathrm{vfov}}$ is the angle of the camera’s Vertical Field of View (VFOV).
III-C Attitude Loop Controller
The desired acceleration is derived from the difference between the desired and current velocities:
$\mathbf{a}_{\mathrm{d}}=\mathbf{V}_{\mathrm{d}}-\mathbf{V}_{\mathrm{now}}$ (17)
where $\mathbf{V}_{\mathrm{now}}$ denotes the current velocity. Based on the multicopter’s dynamics model and the equation ${}^{\mathrm{e}}\mathbf{f}_{\mathrm{d}}=f_{\mathrm{d}}\mathbf{n}_{\mathrm{fd}}$, the normalized direction of the lift force is:
$\mathbf{n}_{\mathrm{fd}}=\frac{\mathbf{a}_{\mathrm{d}}-\mathbf{g}}{\parallel\mathbf{a}_{\mathrm{d}}-\mathbf{g}\parallel}.\textbf{}$ (18)
The attitude $\mathbf{R}_{\mathrm{d}}$, which incorporates only pitch and yaw, is designed to achieve the desired direction of lift. Here, the rotation matrix that includes only pitch and yaw angles is denoted as $\mathbf{R}_{\mathrm{title}}$. Thus, the desired attitude angle can be expressed as:
$\mathbf{R}_{\mathrm{d}}$ $=\mathbf{R}_{\mathrm{tilt}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}},$ (19)
$\mathbf{R}_{\mathrm{tilt}}$ $=\mathbf{I}+[\mathbf{r}]_{\times}\sin\phi+\left[\mathbf{r}\right]_{\times}^{2}\left(1-\cos\phi\right)$
where $\mathbf{r}=\mathbf{n}_{\mathrm{f}}\times\mathbf{n}_{\mathrm{fd}}$, $\phi=\mathrm{arccos}\left(\mathbf{n}_{\mathrm{f}}^{\mathrm{T}}\mathbf{n}_{\mathrm{fd}}\right)$. To achieve the desired pose as described above, a Lyapunov candidate function is formulated as:
$L=\mathrm{tr}\left(\mathbf{I}-\mathbf{R}_{\mathrm{d}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right).$ (20)
Given that $\left\|\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right\|_{\mathrm{F}}=\sqrt{2\mathrm{tr}\left(\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right)}$, $L=\left\|\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right\|_{\mathrm{F}}^{2}/2\geq 0$, the derivative of the candidate function is:
$\dot{L}$ $=-\mathrm{tr}\left(\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}[^{\mathrm{b}}\bm{\omega}]_{\times}\right)$ (21)
$=\mathrm{vex}\left(\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}-\mathbf{R}_{\mathrm{b}}^{\mathrm{eT}}\mathbf{R}_{\mathrm{d}}\right)^{\mathrm{T}}{\mathrm{b}}\bm{\omega}.$
To make $\dot{L}\leqslant 0$ for stability, the pitch and roll angular velocity controllers are designed as:
${{}^{\mathrm{b}}\bm{\omega}_{1}}=-\mathrm{vex}\left(\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}-\mathbf{R}_{\mathrm{b}}^{\mathrm{eT}}\mathbf{R}_{\mathrm{d}}\right).$ (22)
In this context, the attitude control loop for the multi-rotor interceptor can be generalized:
$\begin{cases}{}^{\mathrm{b}}\bm{\omega}_{\mathrm{d}}=\mathrm{sat}\left({{}^{\mathrm{b}}\bm{\omega}_{1}}+{{{}^{\mathrm{b}}\bm{\omega}_{2}},\omega_{\mathrm{m}}}\right)\\ f_{\mathrm{d}}=\mathrm{min}\left(\mathrm{max}\left(\mathbf{n}_{\mathrm{f}}^{\mathrm{T}}\left(\mathbf{a}_{\mathrm{d}}-m\mathbf{g}\right),0\right),f_{\mathrm{m}}\right)\\ \end{cases}$ (23)
[…]
$\mathrm{sat}\left(\bm{\omega},w_{\mathrm{m}}\right)=\left\{\begin{matrix}\bm{\omega},&\parallel\bm{\omega}\parallel\leq w_{\mathrm{m}}\\ \frac{w_{m}}{\parallel\bm{\omega}\parallel}\bm{\omega},&\parallel\bm{\omega}\parallel>w_{\mathrm{m}}\\ \end{matrix}\right..$
where it is considered to affect the convergence speed but not the system’s stability[19].
IV SITL EXPERIMENTS
IV-A Platform introduction and experimental design
In this section, the proposed algorithm is validated and compared using SITL simulation experiments with RflySim [30].These simulations replicate the architecture and algorithms of real flight experiments, ensuring high fidelity and ease of portability. First, a multi-angle static target interception experiment is designed to verify the accuracy of the proposed algorithm. Next, moving target interception experiments are performed and compared with other algorithms. These experiments include targets with three common motion models and initial relative positions in three different directions.
Figure 4: SITL Platform and Simulation Scenario.
IV-B Static Target Intercept Simulation Experiments
Fifty static target interception experiments are performed with targets at various locations. The multicopter starts at the position (0,0,10) meters in the EFCS, while the target positions, as shown in Fig. 5(a), are randomly generated between 15 and 35 meters from the multicopter. The multicopter’s initial velocity is set to be directly forward, and $k_{a}=2$, $K_{y}=K_{z}=3$, $k_{p}=-0.03$, $k_{p}=-0.01$ are set.
Experimental results demonstrate that the proposed algorithm accurately intercepts most targets within the FOV. The target distribution in the image plane before interception is shown in Fig. 5(c). The interception error box and CEP are illustrated in Fig. 5(b) and Fig. 5(d), respectively. Compared to previous works [18] and [19], which reported $\mathrm{CEP}_{\left[18\right]}=0.457$ m and $\mathrm{CEP}_{\left[19\right]}=0.332$ m, the proposed algorithm achieves a $\mathrm{CEP}_{\mathrm{Proposed}}=0.089$ m, representing reductions of 80.5% and 73.2%, respectively. This indicates that the proposed algorithm enhances the accuracy of static target interception from the decimeter level to the centimeter level in the simulation experiments.
Figure 5: Experiment results of static target interception simulation: (a) multicopter trajectories, (b) boxplots of interception error, (c) initial tagert positions in image plane, and (d) interception error distribution.
IV-C Moving Target Intercept Simulation Experiments
This experiment evaluates the interception accuracy of the proposed algorithm across various target maneuvering scenarios. Three maneuvering models were tested: constant velocity (CV), constant acceleration (CA), and sinusoidal maneuver (SM). Parameters were consistent with those used for intercepting stationary targets. The initial relative positions included three interception scenarios: top-to-bottom, flanking, and tailgate, as shown in Fig. 6. Initially, the interceptor multicopter hovers at (0, 0, 10) m. The target follows three different maneuvering models starting from (0, 25, 1) m, (-8, 15, 8) m, and (0, 30, 10) m, respectively. The models are expressed as:
$\begin{cases}\mathrm{CV}:{{}^{\mathrm{e}}\mathbf{v}_{\mathrm{T}}}=\left(0,0,1\right)\\ \mathrm{CA}:{{}^{\mathrm{e}}\mathbf{a}_{\mathrm{T}}}=\left(0.8,0,0.2\right)\\ \mathrm{SM}:{{}^{\mathrm{e}}\mathbf{p}_{\mathrm{T}}}=\left(2\sin\mathrm{(}2\pi/14t),3t,0\right).\\ \end{cases}$
Figure 6: Trajectory of (a) CV, (b) CA, and (c) SM maneuvering model during moving target interception simulation.
Figure 7: Normal velocity and normal acceleration of (a) CV, (b) CA, and (c) SM maneuvering model during moving target interception simulation.
Figure 8: Errors of (a) CV, (b) CA, and (c) SM maneuvering model in moving target interception simulation.
Experimental results demonstrate that the proposed algorithm achieves the highest interception accuracy in the simulations. Fig. 7 show the variations in normal acceleration and velocity during interception. The proposed algorithm maintains low normal acceleration and velocity, which is ideal for minimizing interception errors. Low normal acceleration reduces trajectory deviations, while low normal velocity minimizes the impact of image processing delays on accuracy. Interception errors are illustrated in Fig. 8. Table I confirms that the proposed algorithm outperforms the other two algorithms.
TABLE I: Interception error of experiments.
Experimental indicators PG[22] IBVS [19] Proposed
Errors of CV model (m) 0.71 0.27 0.09
Errors of CA model (m) 0.92 1.76 0.45
Errors of SM model (m) 0.31 0.20 0.04
Additionally, the effectiveness of the FOV holding controller is evaluated by comparing it with another algorithm. The median, quartiles, boundaries, and outliers of the target in the image plane are shown in Fig. 9. The proposed method exhibits a smaller error variation range than the compared algorithm [19], indicating superior field of view holding capability. This demonstrates the proposed algorithm’s greater potential to handle more complex target maneuver scenarios.
Figure 9: Normalized image plane error comparison.
V REAL FLIGHT EXPERIMENTS
V-A Introduction to Flight Experiments
Real flight experiments are designed to verify the robustness and effectiveness of the proposed algorithm under more complex conditions. The algorithms, architectures, and parameter settings are the same as in the simulation: $k_{a}=2$, $K_{y}=K_{z}=3$, $k_{p}=-0.03$, $k_{p}=-0.01$. The real-world experiments are depicted in Fig. 10. The intruder is a red balloon attached to an multicopter by a rope. The interceptor is another multicopter equipped with several components. At the front is a Jetson CSI monocular camera with a 120-degree FOV for capturing images. Behind it is an NVIDIA Jetson Xavier NX, which serves as the onboard computer for processing sensor information and generating control commands. A Pixhawk Nano V5 handles flight control, generating PWM signals to the motors. Additionally, the interceptor is equipped with an RTK module for data analysis target localization (note: this module is not used during the interception process). A video of the flight experiments can be found at https://youtu.be/sQdRCgnRp4g.
Figure 10: Real-world flight experiments and hardware architectures for intercepting drones.
V-B Static Target Intercept Flight Experiment
The interceptor drone starts from a standstill to intercept a red balloon shaking in the wind about 10m away. The results of the experiment are shown in Figure 11. And the trajectory is shown in (a) and (b). As shown in (c), the target always stays in the FOV during interception and moves only a small range. It proves the effectiveness of the proposed FOV holding controller in the real world. The normal acceleration change little during the whole interception process can be seen in (d). And As shown in (e), normal velocity keep at a small level and the maximum velocity is more than 6m/s.
Figure 11: Results of a flight experiment to intercept a static target.
V-C Moving Target Intercept Flight Experiments
Three interception experiments of different maneuver targets are designed: accelerated escape, uniform takeoff, and circular trajectory. These experiments are challenging due to the manoeuvrability of the target and the effects of wind in the field. The relevant initial settings are the same as for the static interception experiment. The interceptor multicopter start the interception task from a standstill. Scenarios and interception trajectory can be seen in Fig. 12. The starting positions are about 10m according to the target.
Figure 12: Schematic of (a) accelerated escape, (b) uniform takeoff, and (c) circular trajectory flight tagets and interceptor trajectories .
All three intercept experiments successfully achieved interception of maneuvering targets. The time taken for interception was approximately 4.8s, 3.9s, and 4.6s, respectively. The velocity and normal acceleration during interception can be seen in Fig. 13. At the end stage of interception, the normal acceleration and normal velocity of the interceptor were essentially the same as those in the hovering stage. The final interception speeds were around 5 m/s. During the interception of the accelerated escape target and the circular trajectory target, a significant portion of the interceptor’s velocity was used for normal velocity in the early stage. This indicates that the PNG works to ensure a flat trajectory by constantly adjusting the relative position before interception. Recorded trajectories further demonstrate the effectiveness of the proposed guidance law in intercepting maneuvering targets.
Figure 13: Normal accelerations, velocities, and normal velocities of the interceptor during the interception of (a) accelerated escape, (b) uniform takeoff, and (c) circular trajectory flight targets.
Additionally, all experiments maintained the target within the FOV during interception. The position of the target in the image plane can be seen in Fig. 14. Generally, all targets moved within a small area in the middle of the image. Despite some dispersion in the first experiment when intercepting the accelerated escape target, the results prove the effectiveness and robustness of the proposed FOV holding controller in intercepting maneuvering targets in real scenarios.
Figure 14: The trajectory of the target in image plane when intercepting a maneuvering target.
V-D Compare Intercept Flight Experiments
However, the two most relevant algorithms [19] and [22] were tested for intercepting an accelerated escape target. This maneuver model is considered the most challenging by previous experiments. As shown in Fig. 15, both algorithms failed during the interception and were subsequently returned by manual remote control. After analysis, the reasons for the interception failure of these two algorithms may be: (1) The overload during interception exceeds the dynamics range of the multicopter, causing the target to exceed the FOV. (2) Random wind disturbances during the experiments result in irregular balloon oscillations. (3) Two algorithms’ interception accuracy are larger than the size of the balloon.
Figure 15: Snapshots of comparative challenging interception experiments for intercepting accelerated escape targets. Both contrasting algorithms of (a) [19] and (b) [22] failed.
VI Conclusion
In this paper, an IBVS algorithm based on PNG was designed to address the problem of intercepting non-cooperative flight targets with a multicopter. The flat trajectory generated by PNG reduces errors due to image processing latency, thereby improving interception accuracy. Additionally, the proposed algorithm demonstrates increased robustness through the FOV holding controller designed for the multicopter. Extensive simulation and real flight experiments have validated the accuracy of the proposed algorithm’s interception capability. However, the proposed algorithm may face challenges when intercepting higher-speed targets. Therefore, future work will focus on enhancing guidance law applications and employing more precise control methods.